OpenStudy (anonymous):

Part A: DO NOT SOLVE! You are only setting up a linear system! (40 points) The table below shows the number of calories burned for a 130-lb person in a minute. If you want to work-out for a total of 60 minutes and burn 500 calories, choose two activities and set up a linear system. Activity Calories Burned per Minute Bicycling 10 Hiking 7 Running 11 Walking 5 Swimming 12 someone please help me

OpenStudy (anonymous):

ok i should learn how to read it says total of 60 minutes and burn 500 calories if you put \(x\) and the number of minutes biking and \(y\) as the number of minutes hiking, then the total being 60 gives and equation of \(x+y=60\) and burning 500 calories tells you \[10x+7y=500\]

OpenStudy (anonymous):

so one system would be \[x+y=60\\10x+7y=500\]

OpenStudy (anonymous):

you can try it again with different activities, get a similar answer one equation will always be \(x+y=60\) the other will change

OpenStudy (anonymous):

SO ONE OF MY EQUATIONS WOULD BE 10X+7Y=500? @satellite73